OpenStudy (anonymous):

3 over m = 12/15 will give a medal

OpenStudy (dmezzullo):

Could u draw this plz?

OpenStudy (anonymous):

\[\frac{ 3 }{ m } = \frac{ 12 }{ 15 }\]You can cross multiply to get: 12m = 45 Then, just divide both sides by 12

OpenStudy (dmezzullo):

sorry thats wat i meant ^^

OpenStudy (anonymous):

ty but i finished studying this i just doing this to give medals

OpenStudy (anonymous):

You'll get:\[m = \frac{ 45 }{ 12 } = \frac{ \cancel3 \times 15 }{ \cancel3 \times 4 }\]